OpenStudy (anonymous):

39 mariposa007 Identify all true statements √9 is a rational # √-1 is NOT a rational # √ab=√a *√b √a+b=√a+√b 9y√3 and 9√3y are like radicals ^3√5 is the symbol for the square root of 5 √a-b ≠ √a-√b

OpenStudy (anonymous):

√-1 is NOT a rational # √ab=√a *√b 9y√3 and 9√3y are like radicals √a-b ≠ √a-√b

OpenStudy (anonymous):

these are the true statements

OpenStudy (anonymous):

this one to ?√a+b=√a+√b